# A Unified Approach to Extract Intepretable Rules from Tree Ensembles via Integer Programming

### v1

@@ Abstract
Tree ensemble methods represent a popular machine learning model, known for their effectiveness in supervised classification and regression tasks. Their performance derives from aggregating predictions of multiple decision trees, which are renowned for their interpretability properties. However, tree ensemble methods do not reliably exhibit interpretable output. Our work aims to extract an optimized list of rules from a trained tree ensemble, providing the user with a condensed, interpretable model that retains most of the predictive power of the full model. Our approach consists of solving a clean and neat set partitioning problem formulated through Integer Programming. The proposed method works with either tabular or time series data, for both classification and regression tasks, and does not require parameter tuning under the most common setting. Through rigorous computational experiments, we offer statistically significant evidence that our method is competitive with other rule extraction methods and effectively handles time series.
Keywords Machine learning Integer programming Interpretability Rule extraction Ensemble learning
@@ 1 Introduction
The need for interpretable and explainable machine learning models has arisen across various applications, leading to a consistently expanding area of research in both the machine learning and operational research communities over the last decade. Explainable Artificial Intelligence (XAI) has emerged as a rapidly growing research field [33], primarily focusing on post-hoc explanation methods to facilitate the understanding of outputs provided by black-box models such as deep neural networks. Although XAI methods can be highly effective in capturing and representing the internal decision-making processes of black-box models, their application in high-risk domains is controversial [51]. Therefore, researchers have gradually shifted their attention to developing inherently interpretable methods.
@@
Commonly regarded as one of the most interpretable machine learning models, decision trees have been for decades the tool of choice for providing highly intuitive and understandable predictions, thanks to their affinity with human reasoning [54]. However, their interpretability is counterbalanced by poor out-of-sample performance, mainly due to overfitting and highly variable predictions. Aggregating trees into ensembles solved this problem, giving rise to models such as random forest [14] or gradient boosted trees [29, 20] that significantly improved performance and robustness for both classification and regression tasks. However, combining a large number of decision trees naturally hinders their original interpretability. In practice, users of tree ensembles can only rely on computing feature importance to attempt to explain the predictions obtained. From a statistical inference viewpoint, feature importance consists of an estimate for the quantity that would have been obtained by making the same prediction with unknown true response function [35]. As extensively discussed in previous research [35, 67, 55], model diagnostics based on feature importance can be highly misleading and biased towards correlated features. For these reasons, our aim is to provide a method that condenses any trained tree ensemble into an interpretable and sparse model while retaining most of its predictive power.
@@
Arguably, rule sets or rule lists represent the only competitor of decision trees in terms of interpretability [65], as they both consist of logical models in the form of if-then statements that are easy to understand [51]. Lists of rules can be built from scratch or by exploiting association rules previously mined from given data [1]. A very interesting connection between decision trees and rule lists involves extracting rules from previously trained tree ensembles. The search space for selecting rules is constrained to be the set of branch nodes of the tree ensemble, which represent the logical conditions that split the training data and make predictions. This approach falls among the post-hoc methods to interpret random forests [3, 27], and it collocates in between the previously discussed areas of research about XAI and interpretable machine learning. Existing methods about extracting rules from tree ensembles involve solving various optimization problems to select the best performing ones, along with their corresponding optimal weight [7, 46, 42, 30]. In most cases, their extension to both classification and regression tasks is not trivial, and they cannot handle more complex data such as time series, images or text. This can be a significant limitation, due to the rising academic and industrial interest in developing machine learning models for complex data.
@@
Recently, Mathematical Programming has played a fundamental role in developing interpretable methods, thanks to its flexibility and ability to induce more sparse and accurate models [17]. A new field belonging to operational research started focusing on formulating Mixed-Integer Linear Programming models to learn optimal classification and regression trees [9, 16, 10], as well as optimal rule lists [61, 39, 62]. The ability to fully control how models learn while minimizing their global prediction error has made this research field growing rapidly, particularly thanks to the recent hardware and algorithmic improvements on solving Integer Programming (IP) problems [9, 39]. However, due to their complexity, most interpretable methods based on Mathematical Programming face scalability issues, both in terms of the number of records and the number of features contained in datasets. In fact, dealing with large-scale data, such as time series, is becoming an increasingly common requirement for machine learning methods related to data-driven decision making [59]. For this reason, developing computationally efficient interpretable methods is a challenging problem. While tree ensemble and rule lists heuristic methods do not suffer from scalability issues, they lack the expressive power and flexibility that Mathematical Programming can offer.
@@
In light of all the above, we present as our contribution a novel method for extracting rule lists from any tree ensemble method by the means of Mathematical Programming. The proposed method relies on solving a well-known and concise IP formulation of a set partitioning problem [5]. The solution of such problem yields the optimal subset of rules in terms of interpretability and predictive power. The use of Mathematical Programming makes the proposed method highly flexible and easily adaptable to various applications. It can be applied to both regression and classification tasks, handling tabular or time series data. Remarkably, the proposed method is agnostic to data structure, and its clean formulation allows users to easily incorporate multiple custom loss functions and robustness measures. Furthermore, under the most common benchmark settings in the literature, the proposed method does not require any parameter tuning. Through rigorous computational experiments, we offer statistically significant evidence that our method is a valid competitor to the state-of-the-art methods for extracting rules from tree ensembles.
@@
The remainder of this paper is organized as follows. In Section 2 we present related works on rule extraction methods from tree ensembles, discussing their main advantages and drawbacks. Moreover, we discuss existing ensemble methods for time series analysis. In Section 3, we provide the mathematical notation and preliminary concept for introducing our problem. In Section 4, we introduce our framework and present the optimization problem. In Section 5, we report our extensive computational results. Finally, Section 6 concludes the paper with a perspective on future works and further research directions.
@@ 2 Related work
Earlier works about learning rule sets from tabular data span multiple fields [39]. The majority of these works focus on classification tasks and do not involve extracting rules from tree ensembles [7]. Various approaches have been explored, including heuristic methods [19, 21] and exact methods based on Integer Programming [61] and Column Generation to reduce the search space [39, 62]. For a comprehensive analysis of rule learning for tabular data, we refer to Fürnkranz et al. [31]. Extracting rules from a pre-trained tree ensemble avoids exploring impractically large search spaces and is considered a state-of-the-art technique for interpreting random forests [3]. Friedman and Popescu [30] proposed a method for extracting lists of rules from both classification and regression tree ensembles, called RuleFit. Their method extracts a subset of rules by minimizing a LASSO optimization problem, selecting the best regularized weight for each rule. However, despite its flexibility, RuleFit tends to perform poorly when extracting rules from deeper trees with highly correlated features [42]. Meinshausen [46] was the first to consider the tradeoff between predictive performance and interpretability with NodeHarvest. The author proposed to extract rules from a tree ensemble by solving a set partitioning problem of non-negative weights through Quadratic Programming. NodeHarvest demonstrates good predictive performance and computational efficiency, offering users an inherently interpretable and sparse model. However, its extension to classification problems is limited to handling no more than two classes. More recently, Bénard et al. proposed a rule extraction method for both classification [7] and regression tasks [6] called SIRUS. Both variants of SIRUS involve a preprocessing step where random forests are restricted to split nodes based on the -empirical quantiles for each feature, with being a user-defined parameter. Then, the extracted rules are filtered based on an user-defined threshold frequency . These filtered rules are then aggregated either by taking their average prediction or by solving a ridge regression problem, depending on the task at hand. Both variants demonstrate good predictive performance and robustness to data perturbation. Despite providing interpretable output, the aforementioned methods are incapable of combining extracted rules by imposing an order on features, resembling a tree-like structure. This issue has been addressed by Liu and Mazumder with FIRE [42], which outperforms aforementioned methods in terms of regression performance and interpretability. The authors proposed a quadratic problem introducing a non-smooth fusion penalty. Their framework promotes solutions that exhibit a tree-like structure, thereby improving interpretability and sparsity of extracted rules. The problem is then solved through an ad-hoc optimization algorithm that leverages its block structure. However, while FIRE excels in regression performance, its extension to classification tasks is not straightforward and has not been investigated by the authors. Moreover, FIRE needs to tune three hyperparameters, requiring computationally expensive validation procedures.
Temporal data.
@@
The previously discussed methods are designed to only handle tabular data. One of the major advantages of the proposed method is its capability to directly handle temporal data. Explainable machine learning methods for temporal data mainly focus on providing users with explanations derived from the data itself, such as time points, subsequences, or instance-based methods [58]. Indeed, the majority of earlier works about knowledge extraction from time series in terms of logical rules involve transitions between time points [59, 43], subsequences [47, 23, 37, 22] or Allen’s temporal logic [53, 2]. Rule discovering methods designed for time series data include various extensions of decision tree algorithms. For instance, Sciavicco and Stan developed a time series extension of the C4.5 algorithm [52], following the work of Brunello et al. [15] who extended ID3. Similarly, previous works introduced tree ensemble methods specifically designed to operate with time series data, such as interval-based boosted classifiers [50] and random forests [26], time points-based random forests [4, 28] and Proximity Forests [44]. Arguably, the most flexible and explainable tree ensemble method for time series is represented by generalized Random Shapelet Forest (gRSF), introduced by Karlsson et al. [38]. The authors were the first to propose a tree ensemble classification and regression method based on shapelets [66, 32], which are generally defined as class-discriminative subsequences of time series. More specifically, these subsequences are the ones that vary the most among different classes of the dataset. The key idea of shapelets-based methods is to extract interesting features by mining local subsequences from time series datasets and utilize them for learning tasks such as classification, regression and clustering. Since shapelets are short time series subsequences, they can be visually inspected, providing users with interpretable output. While the generalized Random Shapelet Forest (gRSF) framework is among the state-of-the-art explainable methods for temporal data, its underlying machine learning model is not interpretable due to its ensemble structure. To the best of our knowledge, our work represents the first rule set extraction method capable of handling tree ensembles operating on time series, thanks to its independence from both data and ensemble structure.
3 Background
@@
In this section, we introduce the main notation and formalism about time series, decision trees and tree ensembles, to prepare the ground for presenting our rule lists learning optimization problem. Our formalism allows us to represent trees and ensemble methods as data-agnostic lists of rules for both classification and regression setting.
@@ 3.1 Time series
A time series is an ordered collection of real values. Given a time series , a time series subsequence of of length starting at position is the ordered collection of values . Please note that there are a total of subsequences. Given two time series , we consider as time series distance the Euclidean norm between the two corresponding vectors . In a more general fashion, distance measures different from the Euclidean norm are allowed. Given two time series and , with , the time series subsequence distance is the minimum distance between and any subsequence of , that is

### v3

# A Unified Approach to Extract Interpretable Rules from Tree Ensembles via Integer Programming

@@ Abstract
Tree ensembles are very popular machine learning models, known for their effectiveness in supervised classification and regression tasks. Their performance derives from aggregating predictions of multiple decision trees, which are renowned for their interpretability properties. However, tree ensemble models do not reliably exhibit interpretable output. Our work aims to extract an optimized list of rules from a trained tree ensemble, providing the user with a condensed, interpretable model that retains most of the predictive power of the full model. Our approach consists of solving a set partitioning problem formulated through Integer Programming. The proposed method works with either tabular or time series data, for both classification and regression tasks, and its flexible formulation can include any arbitrary loss or regularization functions. Our extensive computational experiments offer statistically significant evidence that our method is competitive with other rule extraction methods in terms of predictive performance and fidelity towards the tree ensemble. Moreover, we empirically show that the proposed method effectively extracts interpretable rules from tree ensemble that are designed for time series data.
Keywords Machine learning Integer programming Interpretability Explanation fidelity Rule extraction Tree ensemble Time series classification
@@ 1 Introduction
The demand for interpretable and explainable machine learning models has arisen across various applications. Indeed, decision-makers need to understand the causal mechanisms behind a model’s output in order to justify the choices made, especially in high-risk domains [3]. Consequently, research in interpretability has expanded significantly within both the machine learning and operations research communities over the past decade. Explainable Artificial Intelligence (XAI) has emerged as a rapidly growing research field [45], primarily focusing on post-hoc explanation methods to facilitate the understanding of outputs provided by black-box models such as deep neural networks. Although XAI methods can be highly effective in capturing and representing the internal decision-making processes of black-box models, their application in high-risk domains is controversial [70]. Therefore, researchers have gradually shifted their attention to developing inherently interpretable methods.
@@
Commonly regarded as one of the most interpretable machine learning models, decision trees have been for decades the tool of choice for providing highly intuitive and understandable predictions, thanks to their affinity with human reasoning [74]. However, their interpretability is counterbalanced by poor out-of-sample performance, mainly due to overfitting and inconsistent predictions. Aggregating trees into ensembles solved this problem, giving rise to machine learning models such as random forest [17] or gradient boosted trees [39, 25], which significantly improved performance and robustness for both classification and regression tasks. However, combining a large number of decision trees naturally hinders their original interpretability, producing black-box models. As a consequence, research in the field of interpretable machine learning has focused on developing methods and techniques to interpret both the predictions and the internal mechanisms of tree ensemble models. More in general, methods for interpreting ensemble models such as random forests can be grouped into two categories [3]. The first one refers to internal processing techniques, which involve computing ensemble-specific measures to gain interpretable information about predictions. This category includes various measurement of the feature importance, such as Mean Decrease Accuracy, Mean Decrease Impurity, and Minimal Depth [48, 73]. From a statistical inference viewpoint, feature importance consists of an estimate for the quantity that would have been obtained by making the same prediction with unknown true response function [47]. As discussed in previous research, model diagnostics based on feature importance can be misleading and biased towards correlated features [47, 93, 75]. The second category comprehends all the post-hoc approaches that aim to identify and explain the relationship between tree ensembles and their outputs [3]. These approaches include the so-called born-again tree methods, which aim to induce a single decision tree that replicates the behavior of a given tree ensemble [34, 85, 18]. Another class of post-hoc techniques for interpreting tree ensembles consists of rule extraction methods, which involve inducing a rule-based predictor to capture the predictive power of the ensemble model [34, 3].
@@
Arguably, rule lists (or rule sets) represent the only competitor of decision trees in terms of interpretability [90], as they both consist of logical models in the form of “if-then” statements that are easy to understand [70]. A list of rules can be built from scratch or by exploiting association rules previously mined from given data [1]. An interesting connection between decision trees and rule lists involves extracting rules from previously trained tree ensembles. By doing so, the search space for selecting rules is restricted to the set of branch nodes in the tree ensemble, which represent the logical conditions used to split the training data and make predictions. Thus, extracting rules from a trained tree ensemble avoids exploring impractically large search spaces. Existing methods involve solving various optimization problems to select the best performing rules, along with their corresponding optimal weights [9, 64, 60, 40]. However, these methods lack the flexibility to incorporate diverse loss functions, as they primarily focus on minimizing the squared prediction error. This limitation makes existing approaches difficult to customize and adapt to different learning tasks. Indeed, their extension to both classification and regression tasks is challenging, and their implementations are currently incompatible with complex data types such as time series, images, or text. This can become a significant drawback, especially given the growing academic and industrial interest in developing machine learning models for these types of data. For these reasons, our aim is to provide an explanation method that condenses any trained tree ensemble into an interpretable list of rules, retaining most of its predictive power and faithfully representing its inner structure, regardless of the type of data.
@@
To cope with the lack of flexibility, Mathematical Programming has recently played a fundamental role in developing interpretable and customizable methods, inducing more sparse and accurate models [22]. A new field belonging to operations research started focusing on formulating Mixed-Integer Linear Programming models to learn optimal classification and regression trees [11, 21], as well as optimal rule lists [86, 56, 87]. The ability to fully control how models learn while minimizing their global prediction error has made this research field growing rapidly, particularly thanks to the recent hardware and algorithmic improvements on solving Integer Programming (IP) problems [11, 56]. However, due to their complexity, most interpretable methods based on IP problems face scalability issues, both in terms of the number of records and features contained in data. In fact, dealing with large-scale data is becoming an increasingly common requirement for machine learning methods related to data-driven decision making [80]. For this reason, developing computationally efficient interpretable methods is a challenging problem. While tree ensemble and heuristic rule extraction methods do not suffer from scalability issues, they lack the expressive power and flexibility that Mathematical Programming can offer.
1.1 Our contribution
@@
In light of all the above, we present as our contribution a novel method for extracting an optimal list of rules from any tree ensemble model by the means of Mathematical Programming. The proposed method relies on solving a well-known IP formulation of the set partitioning problem [6]. The solution of such problem yields an optimal subset of rules in terms of interpretability and predictive power. The use of Mathematical Programming makes the proposed method highly flexible and easily adaptable to various applications. It can be applied to both regression and multi-class classification tasks, explicitly handling tabular and time series data. Remarkably, its formulation allows users to easily incorporate multiple custom loss functions and robustness measures. Through extensive computational experiments, we offer statistically significant evidence that our method is a valid competitor to the state-of-the-art methods for extracting rules from tree ensembles. As fidelity and interpretability are essential for producing valuable explanations that accurately describe black-box models [43, 92], we introduce two novel measure to assess explanation fidelity of surrogate models from tree ensembles. Obtained results show that our approach excels in terms of internal fidelity towards the ensemble model, while achieving competitive scores in external fidelity.
Outline
@@
The remainder of this paper is organized as follows. In Section 2, we present an overview on previous works about learning rule lists from data, with a specific focus on related works on rule extraction methods from tree ensembles, discussing their main advantages and drawbacks. Additionally, we discuss the role of explanation fidelity in the context of rule extraction, and we review existing works on explainable methods for time series analysis. In Section 3, we provide the mathematical notation and preliminary concept for introducing the problem, proposing two novel measures of internal fidelity for surrogate models from tree ensembles. In Section 4, we introduce our framework and present the optimization problem, outlining the procedures for validating the proposed method and performing data inference. In Section 5, we report and discuss our computational results in terms of predictive performance and explanation fidelity on classification and regression tasks using benchmark tabular datasets. Furthermore, we conduct a computational evaluation of the proposed method on time series multi-class classification tasks. Finally, Section 6 concludes the paper with a perspective on future works and further research directions.
@@ 2 Related work
In this section, we review and discuss previous works on learning of rule sets from tabular data, with a particular focus on the past literature about rule extraction from ensemble models. Moreover, we discuss the role of explanation fidelity of surrogate models, such as rule sets, with respect to the black-box models from which they are derived. Finally, we explore related work on explainable methods for time series data, aiming to highlight the position of this paper in relation to existing research gaps. Remarkably, earlier works about learning rule sets span multiple fields [41]. Various approaches have been explored, including Disjunctive Normal Form learning theory [81, 53, 38, 76], contrast set learning [5, 7, 54] and frequent pattern mining [36, 42]. The problem of learning rule lists directly from data has been well studied [55, 35], involving Bayesian frameworks [57], heuristic methods [24, 26, 27, 28] and exact methods based on Mathematical Programming [86, 56, 87].
2.1 Rule extraction from tree ensembles
@@
Extracting rules from a trained tree ensemble avoids exploring impractically large search spaces and is considered a state-of-the-art post-hoc approach for interpreting random forests [3, 50]. Friedman and Popescu [40] first proposed a specialized method for extracting lists of rules from both classification and regression tree ensembles, called RuleFit. Their method extracts a subset of rules by minimizing a LASSO optimization problem, selecting the best regularized weight for each rule. However, despite its flexibility, RuleFit tends to perform poorly when extracting rules from deeper trees with highly correlated features [60]. Meinshausen [64] was the first to consider the tradeoff between predictive performance and interpretability with NodeHarvest. The author proposed to extract rules from a tree ensemble by solving a set partitioning problem of non-negative weights through Quadratic Programming. NodeHarvest demonstrates good predictive performance and computational efficiency, offering users an inherently interpretable and sparse model. However, its extension to classification problems is limited to handling no more than two classes. More recently, Bénard et al. proposed a rule extraction method for both classification [9] and regression tasks [8] called SIRUS. Both variants of SIRUS involve a preprocessing step where random forests are restricted to split nodes based on the -empirical quantiles for each feature, with being a user-defined parameter. Then, the extracted rules are filtered based on an user-defined threshold frequency . These filtered rules are then aggregated either by taking their average prediction or by solving a ridge regression problem, depending on the task at hand. Both variants demonstrate good predictive performance and robustness to data perturbation. Nonetheless, its classification variant cannot handle multi-class classification problems. Despite providing interpretable output, the aforementioned methods are incapable of combining extracted rules by imposing an order on features, failing to resemble a tree-like structure. This issue has been addressed by Liu and Mazumder with FIRE [60], which outperforms existing methods in terms of regression performance and interpretability. The authors proposed a quadratic problem introducing a non-smooth fusion penalty. Their framework promotes solutions that exhibit a tree-like structure, thereby improving interpretability and sparsity of extracted rules. The problem is then solved through an ad-hoc optimization algorithm that leverages its block structure. However, while FIRE excels in regression performance, its extension to classification tasks is not straightforward and has not been investigated by the authors. Moreover, FIRE does not offer a direct control on the number of extracted rules.
2.2 Explanation fidelity
In the literature, the discrepancy between a black-box (or opaque) model and a surrogate model extracted from it is commonly referred to as explanation fidelity [65, 83, 50] or completeness [43]. A surrogate model whose predictions closely match those of the opaque model is considered highly faithful. In the context of rule extraction, explanation fidelity requires the extracted rules to faithfully represent the behavior of the trained black-box model. As argued by Zhou [94] and Johansson et al. [50], rule extraction methods should be used to accomplish two distinct tasks. The first one is to understand the underlying reasoning behind individual predictions made by a black-box model, while the second one consists of extracting an interpretable model that serves as a more understandable predictor. Methods handling the former task are evaluated based on fidelity, whereas out-of-sample accuracy is the primary criterion for methods focused on the latter task. Related works on rule extraction methods from tree ensembles fall into the second category. Indeed, the reported results are compared based on predictive performance, while any form of fidelity measure is neglected. According to Gilpin et al. [43], fidelity is necessary to produce explanations that accurately describe the black-box model. In fact, interpretability and fidelity are both needed to obtain valuable explanations [92]. However, they can conflict, as developing an accurate surrogate often requires a more complex model, which can reduce its interpretability [94, 43, 50, 65]. Therefore, a tradeoff between interpretability and fidelity comes into play, and explanation methods should not be evaluated based on a single point along this tradeoff [43]. As first discussed by Messalas et al. [65], two types of fidelity can be identified. The first, known as external fidelity [84], refers to the degree to which the predictions of the surrogate model agree with those of the opaque model. Common measures of external fidelity include calculating the disagreement between the two models, often expressed as the fraction of unexplained variance or through mean squared error. While external fidelity evaluates the consistency between the predictions of the surrogate and the opaque models, it does not guarantee that the surrogate’s decision-making process is faithful to the original one [65]. This aspect is captured by the internal fidelity, which is notably hard to assess [84]. The model-agnostic approach to assessing internal fidelity involves measuring the alignment between the feature importance rankings provided by the surrogate and opaque models. For example, this can be done by sorting the features of a dataset according to their importance as determined by both models, and then computing the rank correlation (e.g., Kendall’s ) between the two orderings [84, 65]. Currently, there is a lack of model-specific methods for assessing the internal fidelity of rule extraction techniques applied to tree ensembles.
2.3 Explainable methods and tree ensembles for temporal data
@@
Explainable machine learning methods for temporal data mainly focus on providing users with explanations derived from the data itself, such as time points, subsequences, or instance-based methods [79]. Indeed, the majority of earlier works about knowledge extraction from temporal data in terms of logical rules involve transitions between time points [80, 61], subsequences [66, 30, 51, 29] or Allen’s temporal logic [72, 2, 12]. Explainable methods designed to discover knowledge from time series include various extensions of the decision tree induction algorithms. For instance, Sciavicco and Stan developed a time series extension of the C4.5 algorithm [71], following the work of Brunello et al. [19] extending the ID3 algorithm. Similarly, previous works introduced tree ensemble models specifically designed to operate with time series, such as interval-based boosted classifiers [69] and random forests [33], time points-based random forests [4, 37] and Proximity Forests [62]. Arguably, the most explainable tree ensemble model for time series is represented by the generalized Random Shapelet Forest (gRSF), introduced by Karlsson et al. [52]. The authors were the first to propose a tree ensemble classification and regression method based on shapelets [91, 44], which are generally defined as class-discriminative subsequences of time series. More specifically, these subsequences are the ones that vary the most among different classes of the dataset. The key idea of shapelets-based methods is to extract interesting features by mining local subsequences from time series datasets and utilize them for learning tasks such as classification, regression and clustering. Since shapelets are short time series subsequences, they can be visually inspected, providing users with interpretable output. While the generalized Random Shapelet Forest (gRSF) framework is among the state-of-the-art explainable methods for temporal data, its underlying machine learning model is not interpretable due to its ensemble structure. To the best of our knowledge, our work represents the first rule set extraction method whose implementation is directly capable of handling tree ensembles designed for temporal data.
3 Preliminaries
@@
In this section, we introduce the main notation and formalism about time series, decision trees and tree ensembles, to prepare the ground for presenting our rules extraction problem. Our formalism allows us to represent trees and ensemble models as feature-agnostic sets of rules for both classification and regression settings. We also present two novel measures of internal fidelity for surrogate models based on rule extraction from tree ensembles.
@@ 3.1 Time series
A time series is an ordered collection of real values. Given a time series , a time series subsequence of of length starting at position is the ordered collection of values . Please note that there are a total of subsequences. Given two time series , we consider as time series distance the Euclidean norm between the two corresponding vectors . In a general setting, distance measures different from the Euclidean norm are allowed. Given two time series and , with , the time series subsequence distance is the minimum distance between and any subsequence of , that is